# GeV-scale QCD Axion

### v1

@@ Abstract
In order to solve the strong CP problem, we study the possibility that the Peccei–Quinn symmetry is broken below the QCD scale. We find that a QCD axion can be above GeV, and may be among the observed $\eta$ resonances. It is immune to quantum gravity corrections. The only fermion that has a $U(1)$ Peccei–Quinn charge is the right-handed up quark. Flavor-changing neutral currents are surprisingly small, except for a possible but not necessary contribution to $D^{0}$$\overline{D}^{0}$ mixing. All accelerator and astrophysical limits can be evaded. The most significant constraint is the mass splitting between $\pi^{\pm}$ and $\pi^{0}$. In a UV completed model, LHC can look for a dijet resonance in $u\bar{u}$, $u\bar{d}$, $d\bar{u}$ channels.
@@ I Introduction
The strong CP problem is considered to be one of the major mysteries in the standard model of particle physics (see e.g., [Hook:2018dlk] for a recent review). The theory of strong interaction, Quantum ChromoDynamics (QCD), can have a topological term in its Lagrangian
@@
with an estimate $d_{n}=1.48(14)(31)\times 10^{-16}\theta~e~{\rm cm}$ from a recent lattice QCD calculation [Liang:2023jfj]. Given the experimental upper limit $d_{n}<1.8\times 10^{-26}~e~{\rm cm}$ [Abel:2020pzs], the angle $\theta$ must be exceedingly small $\theta\lesssim 10^{-10}$ while theoretically it can be as large as $\pi$ in principle. To understand why an angle is so small is the strong CP problem (see, e.g., [benabou2025clearingstrongcpproblem] for a recent discussion why this is indeed a problem).
@@
where $f_{a}$ is the energy scale where the $U(1)_{PQ}$ symmetry is broken and is called the axion decay constant [Peccei:1977hh, Peccei:1977ur]. The axion settles to the minimum of the potential where $\theta$ is dynamically cancelled with $\langle a\rangle=-f_{a}\theta$. However the original proposal by Peccei and Quinn is experimentally excluded by beam dump experiments and search for $J/\psi,\Upsilon\rightarrow a\gamma$ etc as it assumed $f_{a}\approx v$ where $v=\sqrt{2}\langle H^{0}\rangle=246$ GeV is the energy scale of the electroweak symmetry breaking.
“Invisible” axion models were proposed to make the axion hypothesis phenomenologically viable by assuming $f_{a}\gg v$, Since the axion interaction is suppressed by $1/f_{a}$, a large $f_{a}$ can evade experimental constraints. The concrete models are broadly classified as DFSZ axion models [Dine:1981rt, Zhitnitsky:1980tq] if the standard model quarks and leptons have $U(1)_{PQ}$ charges, or KSVZ axion models [Kim:1979if, Shifman:1979if] if $U(1)_{PQ}$ charges are carried by new colored fermions. Astrophysical constraints typically place a very high lower limit on $f_{a}$. For instance, the proto-neutron star core of SN1987a or neutron stars would cool too fast if $f_{a}\lesssim 5\times 10^{8}$ GeV [Raffelt:2006cw, Hamaguchi:2018oqw]. For $f_{a}\gtrsim 10^{11}$ GeV, the misalignment production or cosmic strings can produce axion as a viable dark matter candidate [Bae:2008ue, Benabou:2023npn]. This is an attractive hypothesis that hits two birds (strong CP problem and dark matter) with a single stone (axion). There is an active effort to search for a dark-matter axion experimentally (see e.g. [Giannotti:2024xhx] for a recent review).
Unfortunately, such “invisible” axion models are subject to potentially disastrous quantum-gravity corrections [Kamionkowski:1992mf, Holman:1992us, Barr:1992qq]. It is often estimated that the $U(1)_{PQ}$ global symmetry is explicitly broken by powers in $f_{a}/M_{Pl}$, and a high $f_{a}$ makes the theory more susceptible to such effects. One way to avoid this problem is to obtain $U(1)_{PQ}$ as an accidental low-energy symmetry for a composite axion [Choi:1985cb, Randall:1992ut] (see e.g., [Gherghetta:2025fip, gherghetta2025highqualityaxionexactsusy] for recent concrete models).
@@ II The Main Idea
Here, $\kappa$ is a Yukawa coupling of $\varphi$ to the up quark. The Lagrangian is classically invariant under $U(1)_{PQ}$ but it is anomalous under QCD because of the rotation of $u_{R}$. As a result, we can remove $\theta$ by the $U(1)_{PQ}$ rotation from the Fujikawa measure of the fermion path integral [Fujikawa:1979ay, Fujikawa:1980eg], $\varphi\rightarrow e^{-i\theta}\varphi$, $u_{R}\rightarrow e^{i\theta}u_{R}$, solving the strong CP problem.
@@ III Up Quark Mass
There is no mass term for the up quark in Eq. (4) because it is forbidden by the $U(1)_{PQ}$ symmetry. However the chiral symmetry breaking in QCD induces a linear term in $\varphi$ which makes it acquire a vacuum expectation value producing the up-quark mass. This mechanism is akin to the bosonic technicolor theory of electroweak symmetry breaking with elementary Higgs doublet [Samuel:1990dq].
@@
Here and below we adopt the PDG averages of quark masses defined in the $\overline{\rm MS}$ scheme at the renormalization scale of 2 GeV as reference values [ParticleDataGroup:2024cfk]. Recall that the matrix element $\mu^{3}=\langle\overline{u}_{L}u_{R}\rangle$ also runs in such a way that the product $m_{q}\mu^{3}$ is renormalization-group invariant. Due to the linear term in $\varphi$ it acquires an expectation value
@@
$m_{u}^{0}=$ $\kappa\langle\varphi\rangle=\frac{\kappa^{2}\mu^{3}}{m_{\varphi}^{2}}\,.$ (11)
Note that the lattice determination of quark masses averaged in [ParticleDataGroup:2024cfk] assume the standard model, and the presence of the pion-axion mixing discussed below may change the values. For consistency, we need to allow for an effective up-quark mass subject to the Kaplan–Manohar ambiguity [Kaplan:1986ru]. With the observation that the quark mass matrix $M={\rm diag}(m_{u},m_{d},m_{s})$ and the combination ${\rm det}M^{*}(M^{\dagger})^{-1}$ have the same chiral transformation property, the up-quark mass can have an additional contribution of $O(m_{d}m_{s}/4\pi f_{\pi})$ which is sizable. When we use the PDG value of the up-quark mass $m_{u}=2.2$ MeV, we take the point of view that it applies to the sum of these two contribution,
$m_{u}=m_{u}^{0}+O(1)\frac{m_{d}m_{s}}{4\pi f_{\pi}}=Rm_{u}^{0},$ (12)
where $R$ is the enhancement factor. This ambiguity has been used in the past to argue that the up-quark mass $m_{u}^{0}$ may actually be zero, making the strong CP problem non-existent. That is not what we attempt here. We rather allow for an enhancement of the order of a few. Given this enhancement, and to obtain the required total mass $m_{u}=2.16\pm 0.04$ MeV, we need
@@ V The Axion Mass
It may be puzzling that the axion mass is determined by an external parameter $m_{\varphi}^{2}$ totally unrelated to the QCD scale. Normally we expect the axion mass to vanish in the absence of non-perturbative QCD dynamics and hence should be proportional to powers in $\Lambda_{\rm QCD}$, normally the second power $m_{a}\propto\Lambda_{\rm QCD}^{2}/f_{a}$. It is not the case here. On the other hand if we turn off the QCD anomaly for $U(1)_{A}$, there is an exact Peccei–Quinn symmetry that is spontaneously broken and hence there must be a massless Nambu–Goldstone boson.
To see all this, we need to study the mass matrix of neutral pseudoscalar states including the non-perturbative QCD effects. We use the standard basis including the $U(1)_{A}$ Nambu–Goldstone boson $\eta_{1}$, which is assumed to have a mass term $\frac{1}{2}m_{1}^{2}\eta_{1}^{2}$ from the non-perturbative QCD effects (see [Witten:1979vv, Veneziano:1979ec] for large $N_{c}$ arguments, and [Csaki:2023yas, Kondo:2025njf] for analytic results using supersymmetry),
@@
In the absence of the non-perturbative QCD effect, $m_{1}^{2}=0$ and indeed there is a massless eigenstate. The axion mass remains approximately $m_{\varphi}$ simply because it is the heaviest eigenstate in this mass matrix.
@@ VI Pion Mass Constraint
The mass matrix Eq. (30) suggests that the mixing between the axion and the neutral pion pushes the mass of the neutral pion down relative to that of the charged pion by
$\Delta m_{\pi^{0}}$ $=\frac{\kappa^{2}\mu^{6}}{4f_{\pi}^{2}m_{\pi}m_{\varphi}^{2}}=-0.22~{\rm GeV}^{3}\frac{\kappa^{2}}{m_{\varphi}^{2}}\ .$ (32)
Normally we interpret the observed difference between the charged pion mass $m_{\pi^{\pm}}=139.6$ MeV and the neutral pion mass $m_{\pi^{0}}=135.0$ MeV as the electromagnetic correction, as studied by the lattice calculation, e.g., in [Fodor:2016bgu, Feng:2021zek]. Here we conservatively require that the mass shift to be at most the observed $m_{\pi^{\pm}}$$m_{\pi^{0}}$ splitting of 4.6 MeV. We need
@@
This limit is in a tension but can be reconciled with the requirement Eq. (13) by the enhancement factor $R$, and is the most significant constraint on our proposal  [^1] We thank Tsutomu Yanagida to point out this constraint.. To settle this question, we need lattice calculations combined with chiral perturbation theory allowing for such a mixing in extracting quark masses from pseudoscalar and baryon octet masses. Note also that a higher-order chiral Lagrangian terms such as $\mu_{2}^{2}{\rm Tr}MUMU$ can raise $m_{\pi}^{0}$ relative to $m_{\pi^{\pm}}$ even though it is usually assumed small $\mu_{2}^{2}\approx\mu^{3}/4\pi f_{\pi}$ based on Naive Dimensional Analysis [Georgi:1992dw].
@@ VII Axion Decay
Axion decays rapidly to $3\pi^{0}$
@@
and $\pi^{0}\pi^{+}\pi^{-}$,
$\Gamma(a\rightarrow\pi^{0}\pi^{+}\pi^{-})=\frac{1}{64\pi^{3}}\frac{\kappa^{2}\mu^{6}}{f_{\pi}^{6}}m_{a}=0.055\kappa^{2}m_{a}.$ (35)
Depending on its mass, the axion may decay also into $\pi^{0}\pi^{\pm}K^{\mp}$, $\eta\pi^{0}\pi^{0}$, $\eta\pi^{+}\pi^{-}$, $\pi^{0}K^{+}K^{-}$, $\pi^{0}\eta\eta$, $3\eta$. Clearly the axion is not dark matter because it decays with a lifetime of the order of $10^{-23}$ sec. The usual “light-shining-through-a-wall” search does not work due to its short lifetime.
Compared to these hadronic final states, most experiments searched for $a\rightarrow e^{+}e^{-}$, which is absent in our case, or $a\rightarrow\gamma\gamma$, negligible in our case. Indeed, the axion decay into $\gamma\gamma$ is suppressed by $(\alpha\kappa m_{u}/4\pi m_{\varphi}^{2})^{2}$ as seen in Fig. 1, which is very small when $m_{u}\ll m_{\varphi}$.
@@ VIII Nuclear Force
while no coupling to the neutron. If we allow for the Kaplan–Manohar ambiguity here, the coupling can be smaller by a factor of $R$.
@@ IX FCNC Constraints
Given that the $U(1)_{PQ}$ charge is flavor-dependent, the axion can induce flavor-changing neutral current (FCNC) processes as for variant axion [Bardeen:1986yb] or flaxion [Ema:2016ops].
@@ X Accelerator Constraints
The paper [Jaeckel:2015jla] studied LEP limits on the $a\gamma\gamma$ coupling using an effective operator $\frac{1}{4}g_{a\gamma\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu}$, and excluded the mass range from MeV to 91 GeV down to $g_{a\gamma\gamma}\gtrsim 10^{-3}$$10^{-2}$. Even though the description of this coupling using an effective operator from the triangle anomaly is valid for high $f_{a}$, it would be misleading for our model. In fact, there was an incorrect prediction that $Z\rightarrow\pi^{0}\gamma$ branching fraction is as large as $10^{-3}$ using the triangle anomaly [Jacob:1989pw]. It has been shown that the calculation based on the triangle anomaly is incorrect when photon is far off-shell (including on-shell $Z$) [Deshpande:1990ej, Hikasa:1990xx, Manohar:1990hu, Pham:1990mf, West:1990ta, Young:1990pt, Schroder:1990nr, Chatterjee:1990eg, Maitra:1994pm, Micu:1995sc] (see Fig. 1). For our model, the process is suppressed by $g_{a\gamma\gamma}\sim\frac{\alpha}{2\pi}m_{u}/q^{2}$ and is negligible at LEP energies $q^{2}\geq m_{Z}^{2}$.
Since $\varphi$ does not have Yukawa couplings to charm or bottom quarks, limits on $J/\psi\rightarrow a\gamma$ or $\Upsilon\rightarrow a\gamma$ do not apply. The coupling to the off-shell photon is suppressed as is the case at LEP.
Beam dump experiments such as [Bross:1989mp] look for an axion that decays dominantly into $e^{+}e^{-}$ or $\gamma\gamma$. The GlueX experiment looked for photoproduction of axion-like particle that decays into $\pi^{0}\pi^{+}\pi^{-}$ but the mass limit extends only up to 720 MeV [GlueX:2021myx].

### v2

@@ Abstract
In order to solve the strong CP problem, we study the possibility that the Peccei–Quinn symmetry is broken below the QCD scale. We find that a QCD axion can be above GeV, and may be among the observed $\eta$ resonances. It is immune to quantum gravity corrections. The only fermion that has a $U(1)$ Peccei–Quinn charge is the right-handed up quark. Flavor-changing neutral currents are surprisingly small. All accelerator and astrophysical limits can be evaded. The most significant constraint is the mass splitting between $\pi^{\pm}$ and $\pi^{0}$. In a UV completed model, LHC can look for a heavy quark pair ${\cal D}\overline{\cal D}$ followed by the decay ${\cal D}\rightarrow W+u$ or a single production $qu\rightarrow q\,{\cal U}$ followed by ${\cal U}\rightarrow uZ,uh$. There can be an $O(1)$ contribution $h\rightarrow u\bar{u}\varphi$ in the measurement of $h\rightarrow gg$ or permille effects on the hadronic $Z$ width at a Higgs factory.
@@ I Introduction
The strong CP problem is considered to be one of the major mysteries in the standard model of particle physics (see e.g., [28] for a recent review). The theory of strong interaction, Quantum ChromoDynamics (QCD), can have a topological term in its Lagrangian
@@
with an estimate $d_{n}=1.48(14)(31)\times 10^{-16}\theta~e~{\rm cm}$ from a recent lattice QCD calculation [36]. Given the experimental upper limit $d_{n}<1.8\times 10^{-26}~e~{\rm cm}$ [3], the angle $\theta$ must be exceedingly small $\theta\lesssim 10^{-10}$ while theoretically it can be as large as $\pi$ in principle. To understand why an angle is so small is the strong CP problem (see, e.g., [9] for a recent discussion why this is indeed a problem).
@@
where $f_{a}$ is the energy scale where the $U(1)_{PQ}$ symmetry is broken and is called the axion decay constant [42, 41]. The axion settles to the minimum of the potential where $\theta$ is dynamically cancelled with $\langle a\rangle=-f_{a}\theta$. However the original proposal by Peccei and Quinn is experimentally excluded by beam dump experiments and search for $J/\psi,\Upsilon\rightarrow a\gamma$ etc as it assumed $f_{a}\approx v$ where $v=\langle H^{0}\rangle=174$ GeV is the energy scale of the electroweak symmetry breaking.
“Invisible” axion models were proposed to make the axion hypothesis phenomenologically viable by assuming $f_{a}\gg v$, Since the axion interaction is suppressed by $1/f_{a}$, a large $f_{a}$ can evade experimental constraints. The concrete models are broadly classified as DFSZ axion models [15, 55] if the standard model quarks and leptons have $U(1)_{PQ}$ charges, or KSVZ axion models [33, 48] if $U(1)_{PQ}$ charges are carried by new colored fermions. Astrophysical constraints typically place a very high lower limit on $f_{a}$. For instance, the proto-neutron star core of SN1987a or neutron stars would cool too fast if $f_{a}\lesssim 5\times 10^{8}$ GeV [44, 25]. For $f_{a}\gtrsim 10^{11}$ GeV, the misalignment production or cosmic strings can produce axion as a viable dark matter candidate [5, 8]. This is an attractive hypothesis that hits two birds (strong CP problem and dark matter) with a single stone (axion). There is an active effort to search for a dark-matter axion experimentally (see e.g. [24] for a recent review).
Unfortunately, such “invisible” axion models are subject to potentially disastrous quantum-gravity corrections [31, 27, 7]. It is often estimated that the $U(1)_{PQ}$ global symmetry is explicitly broken by powers in $f_{a}/M_{Pl}$, and a high $f_{a}$ makes the theory more susceptible to such effects. One way to avoid this problem is to obtain $U(1)_{PQ}$ as an accidental low-energy symmetry for a composite axion [12, 45] (see e.g., [23, 22] for recent concrete models).
@@ II The Main Idea
Here, $\kappa$ is a Yukawa coupling of $\varphi$ to the up quark. The Lagrangian is classically invariant under $U(1)_{PQ}$ but it is anomalous under QCD because of the rotation of $u_{R}$. As a result, we can remove $\theta$ by the $U(1)_{PQ}$ rotation from the Fujikawa measure of the fermion path integral [19, 20], $\varphi\rightarrow e^{-i\theta}\varphi$, $u_{R}\rightarrow e^{i\theta}u_{R}$, solving the strong CP problem.
@@ III Up Quark Mass
There is no mass term for the up quark in Eq. (4) because it is forbidden by the $U(1)_{PQ}$ symmetry. However the chiral symmetry breaking in QCD induces a linear term in $\varphi$ which makes it acquire a vacuum expectation value producing the up-quark mass. This mechanism is akin to the bosonic technicolor theory of electroweak symmetry breaking with elementary Higgs doublet [46].
@@
Here and below we adopt the PDG averages of quark masses defined in the $\overline{\rm MS}$ scheme at the renormalization scale of 2 GeV as reference values [40]. Recall that the matrix element $\mu^{3}=\langle\overline{u}_{L}u_{R}\rangle$ also runs in such a way that the product $m_{q}\mu^{3}$ is renormalization-group invariant. Due to the linear term in $\varphi$ it acquires an expectation value
@@
$m_{u}=$ $\kappa\langle\varphi\rangle=\frac{\kappa^{2}\mu^{3}}{m_{\varphi}^{2}}\,.$ (11)
Note that the lattice determination of quark masses averaged in [40] assume the standard model, and the presence of the pion-axion mixing discussed below may change the values. For consistency, we need to allow for an effective up-quark mass subject to the Kaplan–Manohar ambiguity [32]. With the observation that the quark mass matrix $M={\rm diag}(m_{u},m_{d},m_{s})$ and the combination ${\rm det}M^{*}(M^{\dagger})^{-1}$ have the same chiral transformation property, the up-quark mass can have an additional contribution of $O(m_{d}m_{s}/4\pi f_{\pi})$ which is sizable. When we use the PDG value of the up-quark mass $m_{u}=2.2$ MeV, we take the point of view that it applies to the sum of these two contribution,
$m_{u}^{\it eff}=m_{u}+O(1)\frac{m_{d}m_{s}}{4\pi f_{\pi}}=Rm_{u},$ (12)
where $R$ is the enhancement factor. This ambiguity has been used in the past to argue that the up-quark mass $m_{u}$ may actually be zero, making the strong CP problem non-existent. That is not what we attempt here. We rather allow for an enhancement of the order of a few. Given this enhancement, and to obtain the required total mass $m_{u}^{\it eff}=2.16\pm 0.04$ MeV, we need
@@ V The Axion Mass
Note that what we call an axion here is defined as “a new particle that appears in theories that solve the strong CP problem,” not as “a pseudo-Nambu–Goldstone boson that acquires a mass from the QCD chiral anomaly.”
It may be indeed puzzling that the axion mass is determined by an external parameter $m_{\varphi}^{2}$ totally unrelated to the QCD scale. Normally we expect the axion mass to vanish in the absence of non-perturbative QCD dynamics and hence should be proportional to powers in $\Lambda_{\rm QCD}$, normally the second power $m_{a}\propto\Lambda_{\rm QCD}^{2}/f_{a}$. It is not the case here. On the other hand if we turn off the QCD anomaly for $U(1)_{A}$, there is an exact Peccei–Quinn symmetry that is spontaneously broken and hence there must be a massless Nambu–Goldstone boson.
To see all this, we need to study the mass matrix of neutral pseudoscalar states including the non-perturbative QCD effects. We use the standard basis including the $U(1)_{A}$ Nambu–Goldstone boson $\eta_{1}$, which is assumed to have a mass term $\frac{1}{2}m_{1}^{2}\eta_{1}^{2}$ from the non-perturbative QCD effects (see [53, 51] for large $N_{c}$ arguments, and [13, 34] for analytic results using supersymmetry),
@@
In the absence of the chiral anomaly, $m_{1}^{2}=0$ and indeed there is a massless eigenstate. The axion mass remains approximately $m_{\varphi}$ simply because it is the heaviest eigenstate in this mass matrix.
@@ VI Pion Mass Constraint
The mass matrix Eq. (30) suggests that the mixing between the axion and the neutral pion pushes the mass of the neutral pion down by $\Delta m_{\pi^{0}}^{2}=-2\kappa^{2}\mu^{6}/f_{\pi}^{2}m_{\varphi}^{2}=-2\mu^{3}m_{u}/f_{\pi}^{2}$ relative to that of the charged pion. To the leading order in $1/m_{\varphi}$, we find
$m_{\pi^{\pm}}^{2}$ $=\frac{1}{f_{\pi}^{2}}2\mu^{3}(Rm_{u}+m_{d}),$ (32)
$m_{\pi^{0}}^{2}$ $=\frac{1}{f_{\pi}^{2}}2\mu^{3}((R-1)m_{u}+m_{d}).$ (33)
Normally we interpret the observed difference between the charged pion mass $m_{\pi^{\pm}}=139.6$ MeV and the neutral pion mass $m_{\pi^{0}}=135.0$ MeV as the electromagnetic correction, as studied by the lattice calculation, e.g., in [18, 17]. Here we conservatively require that the mass shift to be at most the observed $m_{\pi^{\pm}}$$m_{\pi^{0}}$ splitting of 4.6 MeV. We need
@@
This limit is in a tension but can be reconciled with the requirement Eq. (13) by the enhancement factor $R$, and is the most significant constraint on our proposal  [^1] We thank Tsutomu Yanagida to point out this constraint.. To settle this question, we need lattice calculations combined with chiral perturbation theory allowing for such a mixing in extracting quark masses from pseudoscalar and baryon octet masses. Note also that a higher-order chiral Lagrangian terms such as $\mu_{2}^{2}{\rm Tr}MUMU$ can raise $m_{\pi}^{0}$ relative to $m_{\pi^{\pm}}$ even though it is usually assumed small $\mu_{2}^{2}\approx\mu^{3}/4\pi f_{\pi}$ based on Naive Dimensional Analysis [21].
@@ VII Axion Decay
Axion decays rapidly to $u\bar{u}$,
$\Gamma(a\rightarrow u\bar{u})$ $=\frac{3}{16\pi}\kappa^{2}m_{a}=0.060\kappa^{2}m_{a},$ (35)
if computed perturbatively, or into pions,
@@
$\Gamma(a\rightarrow\pi^{0}\pi^{+}\pi^{-})$ $=\frac{1}{64\pi^{3}}\frac{\kappa^{2}\mu^{6}}{f_{\pi}^{6}}m_{a}=0.055\kappa^{2}m_{a},$ (37)
if computed with the chiral Lagrangian. They cannot be trusted, but provide rough estimates. Depending on its mass, the axion may decay also into $\pi^{0}\pi^{\pm}K^{\mp}$, $\eta\pi^{0}\pi^{0}$, $\eta\pi^{+}\pi^{-}$, $\pi^{0}K^{+}K^{-}$, $\pi^{0}\eta\eta$, $3\eta$. Clearly the axion is not dark matter because it decays with a lifetime of the order of $10^{-23}$ sec. The usual “light-shining-through-a-wall” search does not work due to its short lifetime.
Compared to these hadronic final states, most experiments searched for $a\rightarrow e^{+}e^{-}$, which is absent in our case, or $a\rightarrow\gamma\gamma$, negligible in our case. Indeed, the axion decay into $\gamma\gamma$ is suppressed by $(\alpha\kappa m_{u}/4\pi m_{\varphi}^{2})^{2}$ as seen in Fig. 1, which is very small given $m_{u}\ll m_{\varphi}$.
@@ VIII Nuclear Force
while no coupling to the neutron.
@@ IX FCNC Constraints
Given that the $U(1)_{PQ}$ charge is flavor-dependent, the axion can induce flavor-changing neutral current (FCNC) processes as for variant axion [6] or flaxion [16].
@@ X Accelerator Constraints
The paper [30] studied LEP limits on the $a\gamma\gamma$ coupling using an effective operator $\frac{1}{4}g_{a\gamma\gamma}aF_{\mu\nu}\tilde{F}^{\mu\nu}$, and excluded the mass range from MeV to 91 GeV down to $g_{a\gamma\gamma}\gtrsim 10^{-3}$$10^{-2}$. Even though the description of this coupling using an effective operator from the triangle anomaly is valid for high $f_{a}$, it would be misleading for our model. In fact, there was an incorrect prediction that $Z\rightarrow\pi^{0}\gamma$ branching fraction is as large as $10^{-3}$ using the triangle anomaly [29]. It has been shown that the calculation based on the triangle anomaly is incorrect when photon is far off-shell (including on-shell $Z$) [14, 26, 38, 43, 52, 54, 47, 11, 37, 39] (see Fig. 1). For our model, the process is suppressed by $g_{a\gamma\gamma}\sim\frac{\alpha}{2\pi}m_{u}/q^{2}$ and is negligible at LEP energies $q^{2}\geq m_{Z}^{2}$.
Since $\varphi$ does not have Yukawa couplings to charm or bottom quarks, limits on $J/\psi\rightarrow a\gamma$ or $\Upsilon\rightarrow a\gamma$ do not apply. The coupling to the off-shell photon is suppressed with $q^{2}=m_{\Upsilon}^{2}$ etc as in the case at LEP.
Beam dump experiments such as [10] look for an axion that decays dominantly into $e^{+}e^{-}$ or $\gamma\gamma$. The GlueX experiment looked for photoproduction of axion-like particle that decays into $\pi^{0}\pi^{+}\pi^{-}$ but the mass limit extends only up to 720 MeV [4].